# Data-driven Implementations of Various Generalizations of Balanced Truncation

### v1

@@ 3 Practical Considerations in Implementation
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big{(}2Real(\sigma_{1})I_{m},\cdots,2Real(\sigma_{v})I_{m}\big{)},$ (100)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big{(}2Real(\mu_{1})[I_{p}-t_{w,1}^{*}\mathscr{H}(\mu_{1})R_{\text{X}}^{-1}\mathscr{H}^{*}(\mu_{1})t_{w,1}]^{-1},\cdots,$
$\hskip 56.9055pt2Real(\mu_{w})[I_{p}-t_{w,w}^{*}\mathscr{H}(\mu_{w})R_{\text{X}}^{-1}\mathscr{H}^{*}(\mu_{w})t_{w,w}]^{-1}\big{)},$ (101)
@@
$\mathscr{H}(\mu_{i})$ =$=\begin{cases}\frac{2Real(\sigma_{i})}{\mu_{i}-\sigma_{i}}\big{[}H(\sigma_{i})-H(\mu_{i})\big{]}H^{*}(\sigma_{i})+H(\mu_{i})D^{T}&\text{if }\sigma_{i}\neq\mu_{i}\\ 2Real(\sigma_{i})H^{\prime}(\sigma_{i})H^{*}(\sigma_{i})+H(\mu_{i})D^{T}&\text{if }\sigma_{i}=\mu_{i}\end{cases},$ (103)
@@
4 Numerical Results
We present two numerical examples. The first is an illustrative example designed to facilitate convenient verification of the key theoretical results presented in the paper. The second is a modest-order RLC circuit model, which demonstrates the performance of the algorithms in a practical setting. For the state-space matrices of the RLC circuit, refer to [3]. The MATLAB codes for reproducing the results of both examples are provided in the Appendices. Readers only need to run these codes to replicate the results presented in this section. All necessary accompanying MATLAB functions are also included in the Appendices. All simulations were conducted in MATLAB R2021b on a laptop with a 2 GHz Intel i7 processor and 16 GB of RAM.
4.1 Example 1: Illustrative Example
Consider the following $8^{th}$-order system with $2$ inputs and $2$ outputs:
$A$ $=\begin{bmatrix}-1.7126&0.8101&0.4490&-0.0665&-0.7850&-0.7015&-0.1898&0.5004\\ 0.8101&-2.2155&-0.2692&0.3775&0.9095&0.4206&0.4722&-0.6070\\ 0.4490&-0.2692&-1.2961&-0.0875&-0.1553&0.4107&-0.0627&0.0416\\ -0.0665&0.3775&-0.0875&-0.8279&-0.1431&0.3780&0.1840&0.1695\\ -0.7850&0.9095&-0.1553&-0.1431&-1.9389&-0.6784&-0.6228&1.0619\\ -0.7015&0.4206&0.4107&0.3780&-0.6784&-1.6593&-0.4140&0.2694\\ -0.1898&0.4722&-0.0627&0.1840&-0.6228&-0.4140&-1.3181&0.5562\\ 0.5004&-0.6070&0.0416&0.1695&1.0619&0.2694&0.5562&-1.4047\end{bmatrix},$
$B$ $=\begin{bmatrix}0.5936&0\\ 0&0.4608\\ -0.3270&0\\ 0.4523&0.1748\\ 0.3379&-0.1452\\ 0.2104&0\\ 0.0376&-0.4913\\ -0.3827&0\end{bmatrix},$
$C$ $=\begin{bmatrix}-0.2573&0.2372&-0.1253&0&0&0.1977&-0.1448&-0.5704\\ 0&0.3321&-0.1056&-0.5248&0.1269&-0.4520&-0.7193&-0.5483\end{bmatrix},$
$D$ $=\begin{bmatrix}0.0708&0.0584\\ 0.0382&0.1167\end{bmatrix}.$
This model is stable, minimum phase, positive-real, and bounded-real, allowing all the algorithms presented in this paper to be verified. The interpolation points used for the algorithms are: $(\sigma_{1},\sigma_{2})=(\mu_{1},\mu_{2})=(0.5,0.7)$.
The step responses of the closed-loop system with $4^{th}$-order LQG and $\mathcal{H}_{\infty}$ controllers—designed using DD-LQG-BT and $\mathcal{H}_{\infty}$-BT—are plotted in Figures 1 and 2. The responses match those of the closed-loop system with $8^{th}$-order intrusively designed LQG and $\mathcal{H}_{\infty}$ controllers.
Figure 1: Performance Comparison of the LQG Controller
Figure 2: Performance Comparison of the $\mathcal{H}_{\infty}$ Controller
By running the MATLAB code in the Appendix A, readers can reproduce the output given in Appendix D, verifying most of the theoretical results in the paper.
4.2 Example 2: RLC Circuit
This model is taken from [3]. The performance of the proposed non-intrusive algorithms is evaluated following the same comparison methodology as in [3]. By running the MATLAB code provided in Appendix B, readers can generate plots demonstrating that the non-intrusive implementations successfully capture all dominant Hankel-like singular values.
Figure 3: Performance Comparison of LQG-BT
Figure 4: Performance Comparison of $\mathcal{H}_{\infty}$-BT
Figure 5: Performance Comparison of PR-BT
Figure 6: Performance Comparison of BR-BT
Figure 7: Performance Comparison of SW-BT
Figure 8: Performance Comparison of BST

### v2

@@ 4 Practical Considerations in Implementation
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\sigma_{1})I_{m},\cdots,2\text{Real}(\sigma_{v})I_{m}\big),$ (98)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\mu_{1})[I_{p}-t_{w,1}^{*}\mathscr{H}(\mu_{1})R_{\text{X}}^{-1}\mathscr{H}^{*}(\mu_{1})t_{w,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\text{Real}(\mu_{w})[I_{p}-t_{w,w}^{*}\mathscr{H}(\mu_{w})R_{\text{X}}^{-1}\mathscr{H}^{*}(\mu_{w})t_{w,w}]^{-1}\big),$ (99)
@@
$\mathscr{H}(\mu_{i})$ $=\begin{cases}\frac{2\text{Real}(\sigma_{i})}{\mu_{i}-\sigma_{i}}\big[H(\sigma_{i})-H(\mu_{i})\big]H^{*}(\sigma_{i})+H(\mu_{i})D^{T}&\text{if }\sigma_{i}\neq\mu_{i}\\ 2\text{Real}(\sigma_{i})H^{\prime}(\sigma_{i})H^{*}(\sigma_{i})+H(\mu_{i})D^{T}&\text{if }\sigma_{i}=\mu_{i}\end{cases},$ (101)
@@
5 Numerical Example
This section evaluates the numerical performance of the proposed algorithms using the same $400^{th}$-order RLC circuit model considered in [3]. The MATLAB codes to reproduce the results are provided in [41]. All simulations were performed in MATLAB R2021b on a laptop with a 2 GHz Intel i7 processor and 16 GB of RAM.
Interpolation points for transfer function sampling are computed using formula (78). The imaginary parts $\omega_{i,\sigma}$ of the interpolation points $\sigma_{i}$ consist of 50 logarithmically spaced frequencies between $10^{-1}$ and $10^{4}$. The real parts are determined by setting $\zeta_{i,\sigma}=10^{-5}$. The points $\mu_{i}$ are identical to $\sigma_{i}$. Transfer function samples are obtained numerically from the state-space realization of the RLC model provided in [3]. The quantities $\sqrt{\lambda_{i}(P_{\text{LQG}}Q_{\text{LQG}})}$, $\sqrt{\lambda_{i}(P_{\mathcal{H}_{\infty}}Q_{\mathcal{H}_{\infty}})}$, $\sqrt{\lambda_{i}(P_{\text{PR}}Q_{\text{PR}})}$, $\sqrt{\lambda_{i}(P_{\text{BR}}Q_{\text{BR}})}$, $\sqrt{\lambda_{i}(PQ_{\text{SW}})}$, and $\sqrt{\lambda_{i}(PQ_{\text{S}})}$ are referred to as Hankel-like singular values.
Figures 16 compare the Hankel-like singular values and the relative error $\frac{\|G(s)-\hat{G}(s)\|_{\mathcal{H}_{\infty}}}{\|G(s)\|_{\mathcal{H}_{\infty}}}$. It can be seen that the $25^{th}$-order ROMs generated by intrusive methods (LQG-BT, $\mathcal{H}_{\infty}$-BT, PR-BT, BR-BT, SW-BT, and BST) and their non-intrusive counterparts (DD-LQG-BT, $D$-$\mathcal{H}_{\infty}$-BT, DD-PR-BT, DD-BR-BT, DD-SW-BT, and DD-BST) accurately capture the $20$ most dominant Hankel-like singular values. Moreover, the proposed data-driven algorithms achieve accuracy comparable to that of the intrusive methods for ROMs of orders ranging from $1$ to $25$.
(a) Hankel-like Singular Values Comparison
(b) Relative Error Comparison
Figure 1: Performance Comparison between LQG-BT and $D$-LQG-BT
(a) Hankel-like Singular Values Comparison
(b) Relative Error Comparison
Figure 2: Performance Comparison between $\mathcal{H}_{\infty}$-BT and $D$-$\mathcal{H}_{\infty}$-BT
(a) Hankel-like Singular Values Comparison
(b) Relative Error Comparison
Figure 3: Performance Comparison between PR-BT and $D$-PR-BT
(a) Hankel-like Singular Values Comparison
(b) Relative Error Comparison
Figure 4: Performance Comparison between BR-BT and $D$-BR-BT
(a) Hankel-like Singular Values Comparison
(b) Relative Error Comparison
Figure 5: Performance Comparison between SW-BT and $D$-SW-BT
(a) Hankel-like Singular Values Comparison
(b) Relative Error Comparison
Figure 6: Performance Comparison between BST and $D$-BST